OpenStudy (lukecrayonz):

How many atoms of Oxygen are there in 18g of water?

OpenStudy (xishem):

\[18g(H_2O)*\frac{1mol(H_2O)}{18.015g(H_2O)}=0.99917mol(H_2O)\]\[0.99917mol(H_2O)*\frac{1mol(O)}{1mol(H_2O)}=0.99917mol(O)\]\[0.99917mol(O)*\frac{6.022*10^{23}atoms}{1mol}=6.0atoms(O)\]

OpenStudy (anonymous):

Xishem, you almost had it. All the math save your last step is spot on, but the units must carry. 1 mol will cancel 1 mol, but 6.022x10^23 isn't really going anywhere. You'll have to figure out the exponent.

OpenStudy (xishem):

Pesky exponents. Here you go.\[0.99917mol(O)*\frac{6.022*10^{23}atoms}{1mol}=6.0*10^{23}atoms(O)\]

OpenStudy (anonymous):

They might be pesky, but most professors will nail you on both exponents and unit labels. Regardless of the nature of the math.

OpenStudy (xishem):

I know :P It's easy to forget to write them on this site. I rarely miss them in writing.

OpenStudy (anonymous):

last semester I word processed my first chm hw assignment, after spending 4 hours doing it out long hand. just so that a class mate could have something to study off of.